Brief Exercise 18-12 a1-a2

The following data were taken from the income statements of Mydorf Company.

2020

2019

Sales revenue $6,470,000 $6,220,000
Beginning inventory 920,000 850,000
Purchases 4,422,000 4,583,500
Ending inventory 1,100,000 920,000
Compute for each year the inventory turnover. (Round answers to 1 decimal place, e.g. 1.6.)

2020

2019

Inventory turnover times times

LINK TO TEXT

Compute for each year the days in inventory. (Use 365 days for calculation. Round answers to 1 decimal place, e.g. 1.6.)

2020

2019

Days in inventory days days
Answer #1

Calculate inventory turnover

2020 2019
Inventory turnover 4242000/1010000 = 4.2 4513500/885000 = 5.1

Calculate days in inventory

2020 2019
Days in inventory 365/4.2 = 86.9 days 365/5.1 = 71.6 days